sadie5:

Which expression will you generate if you combine the like terms in 5x + 6x + y + y + 3y + 4 + 6 + 2? A. 5(x + y + 7) B. 5x + y + 2 C. 11x + 5y + 12 D. x + 8y – 12 E. 2x + 4y – 10

sadie5:

what do you think

supie:

Ok I'll add some parenthesis to make it easier to understand \((5x+6x)+(y+y+3y)+(4+6+2)\) So add \(5x+6x=?\\1y+1y+3y=?\\4+6+2=?\)

sadie5:

2x + 7y = -1

supie:

No-

supie:

What is 5x+6x?

sadie5:

11x

supie:

Yes, so now we have \(\color{hotpink}{11x}+\color{aqua}{(y+y+3y)}+\color{lime}{(4+6+2)}\) so now what is \(\color{aqua}{y+y+3y}\)?

supie:

or 1y + 1y + 3y

sadie5:

5y

supie:

Yes, so now add 4+6+2 \(\color{hotpink}{11x}+\color{aqua}{5y}+\color{lime}{(4+6+2)}\)

sadie5:

11x + 5y + 12

supie:

Yes, correct.